The amount of money that Gali has to invest at a simple interest rate of 15% per year, to earn R5 250 interest after three years, is

1.     R5 249,48.

2.     R122 500,00.

3.     R11 666,67.

4.     R10 000,00

5.     R3 620,69.


1
Expert's answer
2021-05-17T16:20:00-0400

"5250 = X \times 0.15 \times 3 \\

5250 = 0.45X \\

X = \frac{5250}{0.45} \\

= 11666,67"

Answer: variant 3. R11 666,67